B Question Hel Elkland Heating & Cooling installs and services commercial heating and cooling systems. Elkland uses job costiData Table Job 101 Job 102 Direct materials used..... $ 19,500 $ 10,500 Direct. labor hours used.... 185 7 2 Print Done
Answer #1

1) Predetermine overhead rate

Estimated overhead / estimated labor hour = Overhead rate
66750 / 4450 = 15 per hour

2) Total overhead cost

Job 101 Job 102
Total overhead cost 185*15 = 2775 72*15 = 1080

3) Job Cost

Job 101 Job 102
Direct material 19500 10500
Direct labor 185*28 = 5180 72*28 = 2016
Overhead 2775 1080
Total job Cost 27455 13596